Consider the following situation.

a right-tailed test with α = 0.1

Find the appropriate rejection regions. (Round your answers to two decimal places. If the test is one-tailed, enter NONE for the unused region.)

z >

z <

State your conclusion if the observed test statistic was z = 2.17. If appropriate, provide a measure of reliability for your conclusion.

a. The null hypothesis is not rejected.

b. The null hypothesis is rejected at the 0.1% level.

c. The null hypothesis is rejected at the 10% level.
Answer #1

Solution,

This is right tailed test,

α = 0.1

P( z > z ) = 0.1

= 1 - P(Z < z) = 0.1  

= P(Z < z) = 1 - 0.1

= P(Z < z ) = 0.9

= P(Z < 1.28 ) = 0.9  

z > 1.28

test statistic = z = 2.17

test statistic > critical value

c. The null hypothesis is rejected at the 10% level.